### v1

# EvidenceMap: Unleashing the Power of Small Language Models with Evidence Analysis for Biomedical Question Answering

Abstract
Current LLM-based approaches improve question answering performance by leveraging the internal reasoning abilities of models or incorporating external knowledge. However, when humans address professional problems, it is essential to explicitly analyze the multifaceted relationships from multiple pieces and diverse sources of evidence to achieve better answers. In this study, we propose a novel generative question answering framework for the biomedical domain, named EvidenceMap, which explicitly learns and incorporates evidence analysis with small language models (SLMs). The framework describes an evidence map for each question and fully utilizes an SLM to derive the representation of the supportive evaluation, the logical correlation, and the summarization of the related evidence, which facilitates an analysis-augmented generation with another SLM in an autoregressive way. Extensive experiments have shown that introducing an evidence analysis learning process can significantly outperform larger models and popular LLM reasoning methods.
EvidenceMap: Unleashing the Power of Small Language Models with Evidence Analysis for Biomedical Question Answering
Chang Zong, Jian Wan, Lei Zhang thanks: Corresponding author. School of Information and Electronic Engineering Zhejiang University of Science and Technology zongchang@zju.edu.cn
1 Introduction
Answering specialized questions often requires reference to a wider range of professional information, such as academic literature and expert knowledge, compared to general domain inquiries Louis et al. (2024); Singhal et al. (2025). Generally, as shown in Figure 1, humans require the integration of evidence from multiple sources to solve professional problems Lu et al. (2022). This process involves a series of analytical steps, including evidence evaluation Watts et al. (2009), evidence correlation Chen et al. (2021), and evidence summarization Hernes et al. (2015), which ultimately leads to a well-reasoned conclusion.
Figure 1: A diagram illustrates the evidence analysis process when addressing professional questions.
To address the generative question answering task supported by multiple pieces of evidence, existing modern methods center on large language models (LLMs), employing implicit reasoning over multiple relevant pieces of knowledge or explicitly generating multiple thinking steps to mitigate model hallucinations and achieve more accurate answer generation, such as the families of chain-of-thought (CoT) Wei et al. (2022); Lightman et al. (2023); Yao et al. (2024) and retrieval augmented generation (RAG) Lewis et al. (2020); Asai et al. (2023); He et al. (2024); Wang et al. (2024), respectively. However, as illustrated in Figure 2, the reasoning frameworks mentioned above overlook the explicit definition of analytical stages and the implicit learning of analytical skills, failing to accurately simulate the human solving process, perpetuating hallucinations and error propagation problems, especially in professional domains.
Figure 2: The paradigm of our framework (right) differs from methods such as CoT and RAG by employing SLMs for learning evidence analysis expicitly, thereby reducing error propagation during reasoning.
To further alleviate the aforementioned issues, this study focuses on biomedical question answering and proposes a novel framework, named EvidenceMap. The framework enables explicit learning and utilization of evidence analysis with small language models (SLMs), thus efficiently simulating the human-solving process. Specifically, our method first acquires multiple pieces of evidence from different sources, and then constructs an evidence map, utilizing a small pre-trained language model (S-PLM) to learn latent representations for evidence evaluation, evidence correlation, and evidence summarization, which serve as the output of the evidence analysis. Finally, these representations, along with textual input, are delivered into a small generative language model (S-GLM) to autoregressively derive answers. The experimental results demonstrate that using SLMs for learning and employing evidence analysis is a promising generative question-answering approach.
Our contributions are summarized as follows.
  • We propose EvidenceMap, a novel framework for generative biomedical question answering, which mimics the analysis of evidence from humans. The framework consists of the acquisition, evaluation, correlation, and summarization of evidence and improves the efficacy of utilizing multiple pieces of evidence.
  • We implement an efficient method for learning evidence analysis in latent space, which requires only updating a small pre-trained language model and subsequently driving another small generative model to achieve better performance in generative question answering.
  • Extensive evaluations demonstrate that our framework for learning explicit evidence analysis can significantly improve the quality of responses compared to larger language models and popular LLM-based reasoning methods.
2 Definitions
2.1 Evidence Map Modeling
We construct an evidence map for each input question to facilitate reasoning during generation. An evidence map introduces three key types of information, including the evaluation of how the evidence supports the question ($\mathcal{R}_{eval}$), the correlation between the evidence ($\mathcal{R}_{cor}$), and the summarization of all the evidence ($E_{sum}$). Generally, an evidence map can be formulated as follows.
$\mathcal{M}$ =$=\{Q,\mathcal{E}_{1},\mathcal{E}_{2},...\mathcal{E}_{m},\mathcal{R}_{eval},\mathcal{R}_{cor},E_{sum}\},$ (1)
$E_{sum}$ $=Agg(\mathcal{E}_{1},\mathcal{E}_{2},...\mathcal{E}_{m}),$
where $Q$ is the question to be addressed, $Agg(\cdot)$ is the aggregation function to derive the summarization $E_{sum}$, $\mathcal{E}_{s}$ represents a set of evidence from the source $s$, which is denoted as follows.
$\mathcal{E}_{s}$ =$=\{E_{s}^{1},E_{s}^{2},...,E_{s}^{n}\},$ (2)
where $E_{s}^{i}$ is the $i$-th evidence from the source $s$. Furthermore, the supportive evaluation and logical correlation of evidence can be described as below.
$R_{eval}^{i}$ $=[E_{s}^{i},Q,\rightarrow],\forall R_{eval}^{i}\in\mathcal{R}_{eval},$ (3)
$R_{cor}^{(i,j)}$ $=[E_{s}^{i},E_{s}^{j},\leftrightarrow],\forall R_{cor}^{(i,j)}\in\mathcal{R}_{cor},$
where the symbol $\rightarrow$ and $\leftrightarrow$ represent the direction of these two relationships, respectively.
2.2 Learning Generative QA with SLM
The objective of our task is to generate the best answers to biomedical questions by learning evidence analysis from the evidence map, which can be described as follows.
$A^{*}=f_{glm}\big{(}Q,f_{plm}(\theta^{*}),\mathcal{M}\big{)},$ (4)
where $f_{plm}(\theta^{*})$ is a tiny pre-trained model (e.g. BERT-base), $f_{glm}$ is a small generative model (e.g. Llama-3.2-3B) for providing the best answer $A^{*}$. Then, our learning process is to find the optimized parameter set $\theta^{*}$, which leads to a minimum generative loss for all question-answer pairs, given the corresponding evidence map $EM_{i}$ for each question $Q_{i}$. This can be denoted as below.
$\theta^{*}=\mathop{\arg\min}\limits_{\theta}\sum_{i}\mathcal{L}\Big{(}A_{i},f_{glm}\big{(}Q_{i},f_{plm}(\theta),\mathcal{M}_{i}\big{)}\Big{)}.$ (5)
Figure 3: The overview of our EvidenceMap framework. An evidence-augmented question is modeled as an evidence map, in which the representations of evidence evaluation, correlation and summarization are computed and updated by a tiny pre-trained model accordingly, in order to facilitate answering by a small generative model.
3 Framework
The EvidenceMap framework is illustrated in Figure 3, in which a pre-trained model is fine-tuned to learn the analysis of evidence to facilitate a small generative model that provides better responses. The three main steps are described below.
3.1 Evidence Acquisition
For each input question, we first gather relevant original evidence from multiple sources, including the intrinsic knowledge of an LLM and the external knowledge of academic papers. Specifically, for the evidence of an LLM, we ask GPT-4o Hurst et al. (2024) to take advantage of its knowledge to provide a series of key points that support the question, described in a concise language (see Appendix A.2). In addition, we extract paper snippets from existing biomedical QA datasets such as BioASQ Krithara et al. (2023) and PubMedQA Jin et al. (2019), which have already been annotated by human experts, as evidence related to a question. Therefore, an evidence-augmented QA dataset is constructed, in which each question corresponds to one piece of LLM evidence $E_{lm}$ and multiple pieces of paper evidence from $E_{p}^{1}$ to $E_{p}^{n}$. The input data can be described as follows.
$Input=\{Q,E_{lm},E_{p}^{1},E_{p}^{2},\cdots,E_{p}^{n}\}.$ (6)
3.2 Latent Evidence Analysis
The Evidence Map.
As defined in Section 2.1, each input question, along with its associated evidence, is modeled as an evidence map. In addition to the evidence pieces themselves, the evidence map also incorporates the analysis among the evidence, including the evaluation of how each piece of evidence supports the question $R_{eval}^{i}$, the correlation between any two pieces of evidence $R_{cor}^{(i,j)}$, and the summarization of all the evidence $E_{sum}$.
SLM-based Evidence Analysis.
We prompt a small pre-trained language model to perform the multifaceted evidence analysis in latent space. Specifically, for evidence summarization, we first obtain the overall semantics of each evidence node, and then use a multi-layer perceptron (MLP) to aggregate and produce a summarized representation of the evidence, which is formulated as below.
$H_{s}^{(1)}$ $=W_{s}^{(1)}\cdot X+b_{s}^{(1)},$ (7)
$E_{sum}$ $=ReLU(H_{s}^{(n)}),$
where $X$ is the concatenated feature of all evidence nodes, denoted as $[X_{lm},X_{p}^{1},\cdots,X_{p}^{n}]$, $H_{s}^{(n)}$ is the embedding of the $n$-th layer. Inspired by PromptBERT Jiang et al. (2022) and PromptEOL Jiang et al. (2024), we prompt the SLM $f_{plm}$ and get the last hidden states of the model as the evidence feature $X_{p}^{i}$, which is described as follows.
$X_{p}^{i}=f_{plm}^{-1}(E_{p}^{i},\texttt{Prompt}_{i};\theta).$ (8)
Meanwhile, the evaluation of how an evidence node supports the question $Q$ is derived from the same SLM by using $\texttt{Prompt}_{e}$ as below.
$R_{eval}^{i}=f_{plm}^{-1}(E_{p}^{i},Q,\texttt{Prompt}_{e};\theta),$ (9)
where $R_{eval}^{i}$ indicates the supportive relationship between the evidence $E_{p}^{i}$ and the question $Q$. Similarly, the correlation between any two evidence nodes $E_{p}^{i}$ and $E_{p}^{j}$ can be formulated as below.
$R_{cor}^{(i,j)}=f_{plm}^{-1}(E_{p}^{i},E_{p}^{j},\texttt{Prompt}_{c};\theta),$ (10)
which indicates the implicit logical relationships among evidence nodes. More details are provided in Appendix A.1.
3.3 Analysis-Augmented Generation
With the result of the analysis of an evidence map, the final answer to the question is provided in an autoregressive way by a generative language model $f_{glm}$. We first convert all textual evidence and the question into an embedding $H_{t}$ using a text embedder, which is the first embedding layer of $f_{glm}$ as follows.
$H_{t}$ $=\text{Embed}([E_{lm},E_{p}^{1},\cdots,E_{p}^{n},Q]).$ (11)
Then the evidence analysis results are projected into the same space with the textual embedding using a linear layer, as a soft prompt as follows.
$\hat{H}_{a}=\hat{W}\cdot[E_{sum},\mathcal{R}_{eval},\mathcal{R}_{cor}]+\hat{b},$ (12)
where $\mathcal{R}_{eval}$ and $\mathcal{R}_{cor}$ are the sets of all the supportive relationships $R_{eval}^{i}$ and logical correlations $R_{cor}^{(i,j)}$, respectively. We then prepend the analysis embedding $\hat{H}_{a}$ to the textual embedding $H_{t}$ as the overall prompt to implement an analysis-augmented generation with the frozen language model $f_{glm}$, which is formulated as below.
$A=f_{glm}(\texttt{Prompt}_{all};\theta^{*}),$ (13)
where $\texttt{Prompt}_{all}=[\hat{H}_{a},H_{t}]$, and $\theta^{*}$ is the frozen parameter set of $f_{glm}$.
3.4 Training Objective
The autoregressive training objective concentrates on training the parameter set of SLM $\theta$ to provide better analysis representations to predict subsequent tokens accurately by the generative language model. We calculate the probability of generating the $i$-th token of the target answer $A$ with the overall prompt and all previous tokens as follows.
$\mathcal{L}=\sum_{i=1}^{L}log\big{(}A_{i}|\texttt{Prompt}_{all},A_{0:i-1};\theta\big{)}.$ (14)
4 Evaluation
4.1 Implementation Details
In the step of learning evidence analysis, we employ DistilBERT Sanh (2019), a lighter but faster model distilled from BERT Devlin et al. (2019), as the SLM to be updated. In the generation step, the 1B and 3B versions of Llama-3.2 Dubey et al. (2024) are used to provide answers. We conduct experiments on two popular biomedical datasets, BioASQ Krithara et al. (2023) and PubMedQA Jin et al. (2019). During training, we set the learning rate as 5e-4, the batch size as 4, and the maximum number of paper evidence per question as 5. For each experiment, we train the pre-trained SLM for 8 epochs in a single NVIDIA A6000 GPU and evaluate the model on test sets.
4.2 Evaluation Metrics
Reference-based quality aims to evaluate the similarity between generated responses and reference answers, encompassing aspects of linguistic structure and semantics. ROUGE-L Lin (2004) and BERTScore (denoted as BERT-S) Zhang et al. (2019) are popular metrics within this category. LLM-based accuracy (denoted as LLM-ACC) addresses the shortcomings of traditional metrics in assessing precision. Based on existing work such as RAGAS Es et al. (2024) and Self-Evaluation Huang et al. (2024), we prompt GPT-4o and also provide reference answers to assess the generated content, balancing accuracy and overall quality. Additional details are provided in Appendix A.3.
4.3 Main Results
In order to demonstrate the effectiveness of learning evidence analysis, we consider three configurations: 1) Inference-LLM: Using larger language models in general and biomedical domains, including BioMed-Llama Pal and Sankarasubbu (2024), Llama-3.1 Dubey et al. (2024), and GPT-4o-mini Hurst et al. (2024). 2) Inference-SLM: Using a smaller frozen generative model with different reasoning strategies, such as RAG Lewis et al. (2020); Gao et al. (2023), CoT Wei et al. (2022) and RAT Wang et al. (2024); 3) Tuning-SLM: Adapting the task by fine-tuning a generative model with LoRA Hu et al. (2021) or updating the pre-trained SLM via a prompting tuning style Lester et al. (2021); Li and Liang (2021). More details about baseline settings are provided in Appendix B.
Setting Method Size BioASQ PubMedQA
ROUGE-L BERT-S LLM-ACC ROUGE-L BERT-S LLM-ACC
4o-mini-RAG - 0.1549 0.5848 0.7339 0.1328 0.5501 0.5874
Inference-LLM Llama3.1-RAG 8B 0.1936 0.6077 0.7401 0.1835 0.6431 0.6440
OpenBioLLM 8B 0.2196 0.5981 0.7655 0.1936 0.5726 0.6276
Llama3.2-RAG 3B 0.2956 0.6225 0.7643 0.1942 0.6102 0.6361
Inference-SLM Analysis-CoT 3B 0.2519 0.6044 0.6231 0.1918 0.6053 0.5935
Llama3.2-RAT 3B 0.2315 0.5996 0.6948 0.1882 0.6061 0.6125
Prefix-Tuning 3B 0.3497 0.6735 0.7062 0.2175 0.6258 0.6182
LoRA-Tuning 3B 0.3914 0.6920 0.7217 0.2517 0.6319 0.6225
Tuning-SLM RAG-LoRA 3B 0.3987 0.6834 0.7440 0.2443 0.6454 0.6342
EvidenceMap 1B 0.4045 0.7075 0.7511 0.2548 0.6513 0.6375
EvidenceMap 3B 0.4529 0.7387 0.7905 0.2931 0.6783 0.6610
Table 1: The comparison demonstrates the effectiveness of EvidenceMap across two biomedical datasets. In the inference settings, it surpasses RAG and CoT augmented methods with both the same and larger sizes of models. Our framework also outperforms all the tuning-based methods, even with a smaller generative model.
The results of the performance comparison are shown in Table 1. We can make the following observations:
First, Learning evidence analysis significantly improves the ability to utilize diverse information. When using the same generative language model (Llama-3.2-3B), EvidenceMap outperforms RAG, CoT and RAT, demonstrating that our approach effectively leverages the relationships among various pieces of evidence and between evidence and the question, thereby maximizing the value of multifaceted information.
Second, Learning evidence analysis unleashes the power of SLMs to efficiently resolve biomedical questions. Compared to larger models, including the RAG method with general LLMs (GPT-4o-mini, Llama-3.1-8B) and biomedical LLM (OpenBioLLM-8B), our approach demonstrates superior performance even with a 1B model. This indicates that evidence analysis can effectively unleash the capabilities of SLMs.
Third, Explicitly learning evidence analysis is better than implicit tuning of language models. We compare three strategies in the setting of model fine-tuning: fine-tuning the small pre-trained model with soft prompt from evidence texts, performance efficient fine-tuning (PEFT) tailored for generative models, and PEFT using prompts derived from multiple pieces of evidence. The results indicate that the performance of these implicit learning strategies is significantly inferior to that of explicit learning of evidence analysis.
4.4 Ablation Study on Evidence Analysis
We further investigate the effect of each part in the evidence analysis on overall performance, including supportive evaluation, logical correlation, and content summarization. We systematically remove one part at a time to observe its impact (the percentage of decline denoted as $\Delta$).
Method BioASQ
BERT-S $\Delta$ LLM-ACC $\Delta$
w/o Eval 0.7181 $\downarrow$ 2.79% 0.7712 $\downarrow$ 2.44%
w/o Cor 0.7072 $\downarrow$ 4.26% 0.7653 $\downarrow$ 3.19%
w/o Sum 0.7239 $\downarrow$ 2.00% 0.7856 $\downarrow$ 0.62%
Method PubMedQA
BERT-S $\Delta$ LLM-ACC $\Delta$
w/o Eval 0.6611 $\downarrow$ 2.54% 0.6537 $\downarrow$ 1.10%
w/o Cor 0.6556 $\downarrow$ 3.34% 0.6518 $\downarrow$ 1.39%
w/o Sum 0.6679 $\downarrow$ 1.53% 0.6589 $\downarrow$ 0.32%
Table 2: Ablation study of three parts within the evidence analysis on two datasets. w/o Eval, w/o Cor, and w/o Sum represent removing the evaluation, correlation, and summarization, respectively.
As shown in Table 2, each part contributes to the overall performance, which is consistent with our hypothesis. In detail, the logical correlation has the most significant impact on overall performance on both datasets, indicating that determining the logical relationship between two pieces of evidence is crucial for providing a high-quality answer. Meanwhile, evaluating how evidence supports the question plays a more significant role than summarizing the evidence, indicating that assessing the relationship between evidence and the question improves effectiveness. Furthermore, incorporating evidence analysis has a greater effect on the BioASQ dataset compared to the PubMedQA dataset in general, showing that the effect of evidence analysis is influenced by specific domains.
4.5 Ablation Study on Answer Generation
We further analyze the impact of each module in the answer generation step. We denote the method that does not incorporate textual evidence embeddings as "w/o TE", and the one that eliminates the input of evidence analysis as "w/o EA". We replace the MLP projector with a simple linear layer and denote it as "w/o Proj". Performance results and variations are illustrated in Table 3. The results indicate that both the textual input and the analysis of the evidence have a significant impact on the final performance, which can significantly improve the quality of answers beyond the foundation of basic language model generation. In contrast, the projector used to align the evidence analysis results with the generative language model is just an auxiliary to the overall performance.
Method BioASQ
BERT-S $\Delta$ LLM-ACC $\Delta$
w/o TE 0.6059 $\downarrow$ 17.98% 0.6482 $\downarrow$ 18.00%
w/o EA 0.6236 $\downarrow$ 15.58% 0.6555 $\downarrow$ 17.08%
w/o Proj 0.7212 $\downarrow$ 2.37% 0.7834 $\downarrow$ 0.90%
Method PubMedQA
BERT-S $\Delta$ LLM-ACC $\Delta$
w/o TE 0.5808 $\downarrow$ 14.37% 0.6054 $\downarrow$ 8.41%
w/o EA 0.6099 $\downarrow$ 10.08% 0.6384 $\downarrow$ 3.42%
w/o Proj 0.6741 $\downarrow$ 0.62% 0.6517 $\downarrow$ 1.41%
Table 3: Ablation study on the step of answer generation on two biomedical datasets. TE, EA, and Proj are modules from the step representing textual embedding, evidence analysis, and projector, respectively.
4.6 Impact of Pre-trained SLM
One advantage of our framework is that it requires training only a small pre-trained language model (typically with fewer than 200 million parameters). We further investigate the impact of various pre-trained language models on overall performance. We selected four representative pre-trained models, including DistilBERT Sanh (2019) used in our EvidenceMap, and three larger models (BERT-Base Devlin et al. (2019), RoBERTa Liu (2019) and ModernBERT Warner et al. (2024)). We can observe from the Table 4 that:
Method BioASQ
BERT-S $\Delta$ LLM-ACC $\Delta$
w/ DistilBERT 0.7387 - 0.7905 -
    - Base 66M
w/ BERT-Base 0.7557 $\uparrow$ 0.7841 $\downarrow$
    - Base 110M
w/ RoBERTa 0.7376 $\downarrow$ 0.7684 $\downarrow$
    - Base 125M
w/ ModernBERT 0.7425 $\uparrow$ 0.7970 $\uparrow$
    - Base 149M
Method PubMedQA
BERT-S $\Delta$ LLM-ACC $\Delta$
w/ DistilBERT 0.6783 - 0.6610 -
    - Base 66M
w/ BERT-Base 0.6672 $\downarrow$ 0.6526 $\downarrow$
    - Base 110M
w/ RoBERTa 0.6614 $\downarrow$ 0.6547 $\downarrow$
    - Base 125M
w/ ModernBERT 0.6859 $\uparrow$ 0.6746 $\uparrow$
    - Base 149M
Table 4: Experiment on the impact of using different small pre-trained language models. DistilBERT, despite having the fewest parameters, demonstrates strong capabilities in learning evidence analysis for question answering.
Despite having the fewest parameters, DistilBERT remarkably achieves strong performance. DistilBERT retrains most of the ability of BERT but is much lighter, which aligns well with our framework, further illustrating the training of a suitable pre-trained SLM can produce better results than just selecting larger models in our framework.
The most powerful pre-trained language model can still achieve the best results. ModernBERT, the latest encoder-only model with a larger parameter size and higher efficiency compared to older models, achieves the best results on both datasets. This indicates that as SLM continues to update and advance, the capabilities of our framework will also be further enhanced.
4.7 Impact of Evidence Input
The amount of textual evidence varies in the datasets we used. BioASQ has an average of 26.60 relevant papers, while PubMedQA has an average of only 4.84. To standardize the experimental settings and reduce computational overhead, as mentioned in Section 4.1, we set the maximum number of paper acquisitions at 5 and add evidence of LLM to enhance diversity. Meanwhile, as described in Section 3.1, we introduce the textual evidence for each question from an LLM to enhance the diversity of input. Therefore, in this relatively complex setting, the impact of the quantity and sources of evidence on performance remains worthy of exploration.
Figure 4: The variation of BERT-S and LLM-ACC with the number of papers as evidence (PN) and the inclusion of LLM evidence (LLM) on two datasets (Left for BioASQ and right for PubMedQA).
Figure 4 is the chart showing the performance changes with the increase in the number of papers and the addition of evidence from an LLM. In general, overall performance increases with increasing number of papers and LLM-based evidence provided. However, for BioASQ, the rate of increase in the LLM-ACC metric is faster than that of BERTScore and there is a decline only when the number of papers is 3 and 4 for the two metrics, respectively. Although the experiments on PubMedQA demonstrate a lower LLM-ACC, its growth rate is still faster compared to BERTScore. The experimental results on these two datasets indicate that our framework can effectively utilize a greater quantity or a richer source of evidence, thus improving the overall quality of responses.
5 Case Study
In this section, we qualitatively analyze the effectiveness of our framework in addressing biomedical questions by presenting specific cases. Figure 5 presents two examples of utilizing evidence in distinct ways to formulate responses. In each case, a question is provided alongside relevant evidence, the standard answer, and the answers generated by RAG and our EvidenceMap.
Case 1 illustrates a scenario of factual questions. We notice that the content of the standard answer is consistent with the evidence presented in the paper, both referencing the drug combination "Pembrolizumab plus lenvatinib", which is more accurate and comprehensive than the LLM evidence. The answer provided by EvidenceMap is incomplete, mentioning only the drug "Pembrolizumab". In contrast, the response generated by the RAG method references "T-VEC", which is an entirely different medication, which results in the dissemination of incorrect information and leads to a decrease in accuracy. This indicates that although our framework may provide incomplete information, analyzing the relationships between pieces of evidence can mitigate the hallucination problem in generative models and help prevent the generation of incorrect answers.
Case 2 illustrates a scenario that addresses an explanatory question. By analyzing the standard answers, we can tell that a more complex analytical process is necessary. Accurately addressing this question requires reliance on information from a piece of paper evidence and the LLM evidence, while simultaneously avoiding certain elements from other papers and the misleading part of the LLM evidence. Specifically, the "IL-12 family" mentioned in the standard answer should be derived from the LLM evidence, which points out the "interleukin-12 cytokine family". Our EvidenceMap accurately captures this information element and presents it in the response. Meanwhile, the standard answer mentions "immunoreaction and promoting endometrial cell proliferation" a concept supported by evidence from the second paper. EvidenceMap precisely identifies this content and incorporates it into the response. In contrast, the answer generated by the RAG method overlooks these key pieces of information and utilizes extraneous content ("T cells") from other papers. This case illustrates that our approach effectively utilizes multiple and diverse sources of evidence by analyzing the relationships between them, thereby providing more accurate answers.
Figure 5: Two cases with distinct difficulties demonstrate that EvidenceMap can more effectively utilize content from multiple and diverse evidence through learning evidence analysis, thereby providing higher quality answers.
6 Related Work
6.1 Long-Form Question Answering
Long-Form Question Answering (LFQA) is a task in natural language processing that aims to generate detailed, coherent, and informative answers to address open-ended questions posed by users. Traditional research on LFQA focuses primarily on the method of generative models. ELI5 Fan et al. (2019) proposes an abstract generation model that integrates multi-task objective training. A following work Krishna et al. (2021) designs a system based on sparse attention and contrastive learning to address poor relevance issues. Recent studies concentrate more on retrieval-augmented long-form generation. One work Su et al. (2022) enhances the model with fine-grained key information relevant to the answers to improve the accuracy. WebCPM Qin et al. (2023) focuses on retrieval and synthesis, proposing an interactive web-based search that enables real-time interaction with search engines. Furthermore, evaluating the quality of LFQA is an important area. A study on LFQA evaluation Xu et al. (2023) employs a multidimensional assessment approach, focusing on aspects including factual, accuracy, and coherence. ASQA Stelmakh et al. (2022) proposes a reliable metric that can provide clear correctness standards to address the ambiguous definition of answer quality. Compared to previous studies, our research primarily focuses on the scenario of multiple and diverse evidence in biomedical domain, aiming to efficiently address LFQA tasks by enabling SLMs to explicitly learn evidence analysis.
6.2 Textual Knowledge Augmentation
Textual evidence, as an external knowledge, can significantly enhance the quality of language model generation in professional domains. Methods for knowledge augmentation can be categorized into two types. The first is the augmentation during inference. RAG Lewis et al. (2020), Self-RAG Asai et al. (2023), and FLARE Jiang et al. (2023) belong to this category, which improve generation effectiveness by retrieving relevant documents as factual knowledge and adjusting these texts during inference. Meanwhile, a model can reduce hallucinations from generation by incorporating reasoning knowledge, as demonstrated by CoT Wei et al. (2022), ToT Yao et al. (2024), and CoK Wang et al. (2023a), a series of sequential or structured reasoning steps are generated and executed explicitly. Furthermore, methods with an augmentation mixture of factual and reasoning knowledge can further combine the strengths of both, such as RAT Wang et al. (2024). Another type of augmentation is by tuning the model with domain text. Soft prompt tuning Lester et al. (2021); Yang et al. (2023) uses implicit learnable vectors to replace explicit tokens as prompts to flexibly adapt to multiple downstream tasks. Instruction tuning Wang et al. (2023b); Lin et al. (2023) improves models in knowledge-intensive tasks by constructing specific instruction datasets and learning the utilization of input information and contextual understanding. Compared to the aforementioned methods, our study proposes a novel knowledge augmentation process that prompts the inference of generative models by training the explicit evidence analysis abilities of a tiny pre-trained model.
7 Conclusion
In this work, we present a novel framework for generative biomedical question answering named EvidenceMap, which improves the ability of models to handle multiple and diverse evidence by explicitly training small language models in evidence analysis. The experimental results and cases show that EvidenceMap significantly improves the performance of the generative biomedical question answering by efficient utilization of evidence.
Limitations
We faithfully discuss the limitations that we would like to improve in future work.
First, the current evaluation of our method focuses solely on public datasets in the biomedical domain. Future evaluations will continue to examine its performance in other biomedical datasets, while also conducting tests in additional professional domains.
Second, this study has only tested a limited number of generative language models from the Llama 3 series, and further testing is needed on additional small generative models around the 3B scale, such as Phi-3.5-mini Abdin et al. (2024) and Qwen2.5-3B Yang et al. (2024).
Third, this work focuses primarily on the capabilities of SLMs while incorporating explicit evidence analysis. We will further explore the effects of learning evidence analysis on larger-scale pre-trained and generative models.
Fourth, the generalizability of learning evidence analysis skills and their potential transferability to other models remains an area for further exploration.
Ethical Considerations
Our EvidenceMap framework is not flawless due to limited model capabilities and the possibility of errors or outdated evidence. We test it on two public datasets, BioASQ and PubMedQA, and acknowledge that its effectiveness may be restricted to similar datasets or domains. Its performance in other scenarios is uncertain and poses potential risks. Therefore, it is crucial to exercise caution and verify the accuracy of the answers generated by the method.
Appendix A SLM and LLM Prompts
A.1 Prompts for Evidence Map
The $\texttt{prompt}_{i}$ to obtain the representation of an individual evidence $X_{p}^{i}$ from an SLM:
Evidence: $E_{p}^{i}$ This evidence means:
The $\texttt{prompt}_{e}$ to obtain the representation of the supportive evaluation of evidence $R_{eval}^{i}$ from an SLM:
Evidence: $E_{p}^{i}$ Question: $Q$ How much the evidence supports the question:
The $\texttt{prompt}_{c}$ to obtain the representation of the logical correlation of two pieces of evidence $R_{cor}^{(i,j)}$ from an SLM:
Evidence1: $E_{p}^{i}$ Evidence2: $E_{p}^{j}$ The logical relationship between these two pieces of evidence:
A.2 Prompt for Acquiring LLM Evidence
To simulate the sources of evidence in real problem-solving scenarios, we introduced evidence texts based on the intrinsic knowledge of LLMs, thereby enriching the evidence content. This evidence is analyzed alongside the paper evidence to explore its impact on performance. We prompt GPT-4o to utilize its intrinsic knowledge to acquire evidence for each question as follows.
System Input: You are an AI assistant that helps a human analyst discover evidence that supports the question. User Input: Provide a concise summary that supports answering the question with your own knowledge. The summary should be evidence including key insights that can explain your answer to the question. Question: $Q$ Summary:
A.3 Prompt for LLM-based Metric
We prompt GPT-4o to provide both the accuracy score and the fluency score ranging from 0.0 to 1.0 as follows.
System Input: You are an AI assistant tasked with evaluating the quality of the generated answers in terms of accuracy and fluency. User Input: Please score the generated answer based on accuracy and fluency separately, comparing with the reference answer. The scores should be a float number between 0.0 and 1.0, where 0.0 indicates the generated answer is completely incorrect or unable to understand, 1.0 means the generated answer is totally precise and fluent. Your output should be a dictionary such as {"accuracy": 0.8, "fluency": 0.9}, without any additional information. The question, generated answer, and reference answer are give as below: Question: $Q$ Generated answer: $A$ Reference answer: $A^{*}$ Your output:
Appendix B Details of Baseline Settings
We implement baseline methods by applying various knowledge augmentation strategies based on different pre-trained models for comparison with our framework. For the Inference-LLM setting, we provide evidence the same as EvidenceMap for each question to perform inference on GPT-4o-mini and Llama3.1-8B, implementing a naive RAG approach. We also directly evaluate the adapted OpenBioLLM-8B, which have been tuned by domain datasets. For the Inference-SLM setting, we implement a CoT process based on Llama3.2-3B, providing the three phases of evidence analysis in natural language for the model to follow. In addition, a RAT framework is implemented, enabling the model to generate an initial CoT and optimize the thinking process based on acquired evidence and ultimately perform inference based on the adjusted reasoning process. For the Tuning-SLM setting, we perform prefix-tuning by providing the Llama3.2 model with a soft prompt, which can be optimized with the pre-trained BERT-like model. We also achieve instruction-tuning strategy by constructing evidence-augmented instructions as the dataset, and update Llama3.2-3B using LoRA, which is denoted as RAG-LoRA.
Appendix C Details of Biomedical Datasets
The number of samples and the average amount of evidence per sample in the training and test sets from the two public biomedical datasets are shown in Table 5.
Dataset Set Samples Evidence per Sample
BioASQ Train 5049 12.91
Test 340 26.60
PubMedQA Train 953 4.35
Test 47 4.84
Table 5: Statistical data from the two datasets reveal obvious disparities in the number of samples and average amount of evidence.
Appendix D Results of Language Fluency
We explore the fluency of language generated by our framework. As stated in Appendix A.3, an LLM is employed to score accuracy and fluency of the generated output at the same time. The average values of all samples from the test set are calculated and are shown in Table 6. The results indicate that our framework achieves the best and competitive fluency quality on BioASQ and PubMedQA, respectively.
Method Size BioASQ PubMedQA
LLM-FLU LLM-FLU
4o-mini-RAG - 0.8795 0.8323
Llama3.1-RAG 8B 0.8738 0.8574
OpenBioLLM 8B 0.8662 0.8643
Llama3.2-RAG 3B 0.8910 0.8819
Analysis-CoT 3B 0.8547 0.8541
Llama3.2-RAT 3B 0.8719 0.8675
Prefix-Tuning 3B 0.8592 0.8245
LoRA-Tuning 3B 0.8558 0.8374
RAG-LoRA 3B 0.8734 0.8422
EvidenceMap 1B 0.8829 0.8451
EvidenceMap 3B 0.9088 0.8344
Table 6: The comparison of language fluency measured by LLM on two datasets. Our SLM training architecture saves generation quality on fluency comparing with other language model based methods.